The force needed to keep an object in circular motion is given by F= mv2/r. Combine this with F= BQv to find an expression for the magnetic flux density B needed to keep a charged particle in circular motion for a given radius r and orbital speed v.
1
Expert's answer
2020-05-01T10:13:19-0400

"F=\frac{mv^2}{R}" and "F=qvB"


"\frac{mv^2}{R}=qvB"


"B=\frac{v}{R(q/m)}."